The equilibrium constant (K) is 0.13 at a particular temperature for the reaction: N204(g) 2NO2(9) Given the following sets of initial conditions, what is the net change that must occur for the reaction to reach equilibrium? Does the reaction shift left to reach equilibrium, does the reaction shift right to reach equilibrium or is the reaction at equilibrium at these initial concentrations so no net change will occur? PNo2-0.107 atm, PN20-0.146 atm Po2 -0.206 atm, Po, 0.136 atm PNo2-0.093 atm,P00.066 atm PNo20.056 atm, PN204 - 0.058 atm PNO2-0.106 atm, PN204 -0.052 atm
Answer #1

1)

Qp = p(NO2)^2 / p(N2O4)

= 0.107^2 / 0.146

= 0.0784

since Qp is less than Kp, reaction shift to the right

Answer: shift right

2)

Qp = p(NO2)^2 / p(N2O4)

= 0.206^2 / 0.136

= 0.312

since Qp is more than Kp, reaction shift to the left

Answer: shift left

3)

Qp = p(NO2)^2 / p(N2O4)

= 0.093^2 / 0.066

= 0.13

since Qp is equal to Kp, no change will occur

Answer: no net change

4)

Qp = p(NO2)^2 / p(N2O4)

= 0.056^2 / 0.058

= 0.054

since Qp is less than Kp, reaction shift to the right

Answer: shift right

5)

Qp = p(NO2)^2 / p(N2O4)

= 0.106^2 / 0.052

= 0.216

since Qp is more than Kp, reaction shift to the left

Answer: shift left